# Words2Contact: Identifying Support Contacts from Verbal Instructions Using Foundation Models

### v1

@@ II Related work
II-A Multi-contact whole-body Control
Whole-body controllers are typically formalized as a QP problem with the minimization of a weighted sum of quadratic functions each time-step $t$ [escande2014hierarchical], under several constraints:
$q^{*}(t)$ $=\arg\min_{q(t)}\sum_{i}w_{i}\|c_{i}(t)\|^{2}$
s.t. $\begin{cases}\text{dynamics/kinematics}\\ \text{joint, torque, velocity limits}\\ \end{cases}$
where each $c_{i}(t)$ is a cost function $i$ at time-step $t$ (often called a task), for instance the Euclidean distance between the left hand and its desired position, the distance of the center of mass to the desired position, the head’s orientation etc.
Depending on the robot’s control formulation, the optimization variables $q$ can be the joints’ torque, velocity, or angular position. When multiple contacts are involved, torque control is ideal to directly set the interaction forces and wrenches [henze2016passivity]. Nevertheless, torque control requires both a torque-controlled robot, and an accurate dynamic model. Impedance control can be an alternative to distribute the forces among contacts, but it requires modeling the actuators and lot of parameter tuning [caron2019stair].
In our recent work, we introduced a unified formulation for whole-body multi-contact control [rouxel2024multi], for position-controlled robots to distribute and control the contact forces. The core idea is to utilize the flexibility that stems both the mechanical bending of the robot and the non-perfect joint position controllers. By modeling this flexibility, the joint position commands and the contact forces can be connected, allowing the controller to regulate these forces by incorporating them in the quadratic programming formulation. In this paper, we use this multi-contact whole-body controller to perform experiments with the Talos humanoid robot [stasse2017talos].
@@ II-B Robotics and Language Models
Instead of relying on LLMs solely trained on language, an alternative idea is to use the same learning architectures as LLMs (transformers), but train them on multi-modal robotics data instead of pure text, like in the Robotics Transformers (RT) line of work [brohan2023rt]. A more popular and less compute-intensive approach is to use similar large-scale robotics datasets and incorporate pre-trained language and vision models with a few trained layers to connect the components. OpenVLA [kim2024openvlaopensourcevisionlanguageactionmodel] uses this approach with small open-source models (7-billion parameters), highlighting the potential for substantial achievements with smaller models.
@@ III-A Prompting LLMs
For readers unfamiliar with LLMs, we want to stress the difference between the system prompt and the user prompt. The system prompt is an instruction or message given to the LLM to guide its responses. This prompt sets the tone, context, and boundaries for the conversation, helping the model understand its role and what is expected of it. Considering the constraints of the task, each module has its own system prompt for the LLM. Examples of system prompts are shown in Fig. 3. The user prompt, instead, is the input or query provided by the user to the LLM in natural language.
Figure 3: Some examples of system prompts that are utilized by the LLM in our modules. “+5 examples” refers to the 5 examples that are added to the system prompt, as part of the few-shot prompting technique. All the system prompts are available at: https://hucebot.github.io/words2contact_website/.
@@ III-E Control Module
Once the user confirms that they are satisfied with the current contact point shown to them in image space, we query the LLM one final time, with the end-effector selector prompt (Fig. 3-c), to select the robot’s end-effector that will be used for contact (e.g., right or left hand) and the task type (e.g., support contact or reaching). Then, we extract the 3D position, in camera frame $\mathbf{p}^{cam}=[x,y,z]^{cam}$, from the point cloud.
The Control Module then relies on the SEIKO Retargeting [seiko, seiko2] and Controller [rouxel2024multi] to figure out the commanded motion and smoothly establish the new contact.
The SEIKO Controller takes as input the discrete selection of the end-effector (EE), as well as the 3D position of the selected contact point $[i,j]$ in camera frame $\mathbf{p}^{cam}=[x,y,z]^{cam}$. It transforms it into a position in the robot’s world frame: $\mathbf{p}^{rob}=\mathbf{o}^{rob}_{cam}+\mathbf{R}^{r}_{cam}\mathbf{p}^{cam}$. A spline-based Cartesian trajectory, starting from the current position $\mathbf{x}^{EE}_{t}$ at time $t$, brings the end-effector EE to the desired contact point $\mathbf{p}^{rob}$. SEIKO Retargeting uses a model-based Sequential Quadratic Programming (SQP) optimization to compute feasible whole-body configurations (joint positions and contact forces) that track the Cartesian effector pose commands. The contact forces are automatically determined by the optimization problem of SEIKO, so only the contact locations are necessary, to enforce safety and balance constraints
SEIKO Controller integrates an explicit model of joint flexibility and employs an SQP whole-body admittance formulation to regulate contact forces on our position-controlled humanoid robot. This controller is crucial for performing the distant reaching tasks that challenge the system’s balance, as it enhances robustness against model errors.
@@ IV-B Pilot Study - Evaluation of the correction mechanism and usability of the pipeline
In the first phase, participants were presented with a set of $10$ images, randomly sampled from the same dataset as in Sec. IV-A. Each image displayed a random target marked with a circle of 18 pixels radius; the participant’s task was to tell the system how to accurately place a point, marked with another circle of 5 pixels radius, on the designated target using a maximum of 5 prompts. Each participant received minimal instructions and, notably, was not informed about the existence of the correction module.
In the second phase, we provided participants with an explanation of the system’s functionalities, including two examples of prediction sentences and two correction sentences. With this new knowledge, they instructed the system to identify $10$ targets on $10$ different images.
We measured the distance between the predicted point and the target across prompts, as the distance between the centers of the two circles. The results (Fig. 8) demonstrate a significant improvement in task performance in both phases: all the users quickly learned how to use the system, and were able to bring the point close to the target with few corrections. Their performance was, as expected, better after the prompt expert suggestions (the median distance at the 5th prompt is 21 pixels, which is comparable with the target circle radius). The ability to accurately reach the target in a few iterations confirmed that the prediction-correction mechanism of Words2Contact is efficient and adequate to be used for precise contact identification, which is critical for using this system with real robots. Furthermore, participants reported engagement and satisfaction in using the system.
@@ V Conclusions & Future Work
We introduce Words2Contact, a pipeline for language-guided multi-contact placement for humanoid robots that leverages VLMs and LLMs to interpret verbal commands: this is a powerful tool for teleoperated and collaborative robots. Words2Contact is effective even with small open-source models without fine-tuning. In a pilot study, we show that humans quickly learn to use the system; real robot experiments show that the system provides satisfactory contacts even in difficult environments, thanks to the iterative correction mechanism.
Ongoing work is focused on evaluating the impact of the prediction errors on downstream tasks. Future work will improve the contact prediction using insights from Visual Question Answering (VQA) to better handle abstract spatial concepts (“a bit, a little more”) and link corrections to physical quantities. Additionally, we plan to inform the Contact Prediction with the robot’s dynamics model:

### v2

@@ II Related work
II-A Multi-Contact and Whole-Body Control
Multi-contact tasks present both theoretical and practical challenges that must be addressed carefully. In these scenarios, humanoid robots exhibit redundancies in their whole-body kinematic and contact force distribution [seiko], for example a robot can place its hand on a table in multiple positions, with various combinations of force distribution for each configuration. For humanoid robots, maintaining balance and ensuring the feasibility of commanded motions are crucial for safety, and require consideration of the system’s physical constraints. Robustness is essential for real hardware deployment, necessitating real-time control of posture and contact forces.
Low-level control in humanoid robots is often achieved via Quadratic Programming (QP) for whole-body optimization [escande2014hierarchical], allowing fast computation while accounting for constraints and allowing direct control of contact forces on torque-controlled robots [henze2016passivity].
However, such robots are sensitive to model errors, affecting robustness. In this work, we conducted hardware experiments on the full-size humanoid robot Talos [stasse2017talos] controlled in position. In this context, contact forces can be indirectly tracked through admittance schemes [caron2019stair].
Our proposed contact selection method is agnostic to how multi-contact motion is achieved. We utilized our SEIKO (Sequential Equilibrium Inverse Kinematic Optimization) framework [rouxel2024multi] for the multi-contact experiments. SEIKO uses a Sequential QP to formulate a whole-body admittance controller, explicitly modeling joint flexibility to indirectly regulate contact forces on position-controlled robots.
@@ II-B Robotics and Language Models
Instead of relying on LLMs solely trained on language, an alternative idea is to use the same learning architectures as LLMs (transformers), but train them on multi-modal robotics data instead of pure text, like in the Robotics Transformers (RT) line of work [brohan2023rt]. A more popular and less compute-intensive approach is to use similar large-scale robotics datasets and incorporate pre-trained language and vision models with a few trained layers to connect the components. OpenVLA [kim2024openvlaopensourcevisionlanguageactionmodel] uses this approach with small open-source models (7 billion parameters compared to GPT-3’s 175 billion.), highlighting the potential for substantial achievements with smaller models.
@@ III-A Prompting LLMs
For readers unfamiliar with LLMs, it is important to differentiate between the system prompt and the user prompt. The system prompt is an instruction that guides the LLM’s responses, setting the tone, context, and boundaries for the conversation. Each module has a specific system prompt tailored to its task, as illustrated in Fig. 3. In contrast, the user prompt is the natural language input or query provided by the user.
Figure 3: Some examples of system prompts that are utilized by the LLM in our modules. “+5 examples” refers to the 5 examples that are added to the system prompt, as part of the few-shot prompting technique. All the system prompts are available at: %****␣main.tex␣Line␣525␣****https://hucebot.github.io/words2contact_website/.
@@ III-E Control Module
Once the user confirms their satisfaction with the displayed contact point, we query the LLM one final time using the end-effector selector prompt (Fig. 3-c), to determine the robot’s end-effector (e.g., right or left hand) and the task type (e.g., support contact or reaching). We extract the 3D position, in camera frame $\mathbf{p}^{cam}=[x,y,z]^{cam}$, from the point cloud and we transform it to robot’s world frame using forward kinematics. A spline-based Cartesian trajectory, starting from the current position $\mathbf{x}^{EE}_{t}$ at time $t$, brings the end-effector EE to the desired contact or reaching point $\mathbf{p}^{rob}$.
The Control Module utilizes SEIKO Retargeting [seiko, seiko2] and Controller [rouxel2024multi] to track the effector position, smoothly establishing new contacts by redistributing contact forces across the effector to participate in the robot’s balance. The SEIKO pipeline takes as input the instantaneous effector target position sampled from the spline, along with the contact force measurements from force-torque sensors, and produces joint positions that are sent to the robot’s hardware. SEIKO is implemented as a model-based Sequential Quadratic Programming (SQP) optimization method to initially compute the desired whole-body configuration based on the target effector position. The Controller then regulates contact forces to achieve multi-contact motion while ensuring robustness. For safety, SEIKO enforces motion feasibility and imposes constraints on joint position and velocity limits, quasi-static balance, and conditions to prevent slipping and tipping. If the pipeline receives an infeasible or unreachable position, the robot will attempt to approach the target as closely as possible without compromising its balance.
@@ IV-B Pilot Study - Evaluation of the correction mechanism and usability of the pipeline
In the first phase, participants were presented with a set of $10$ images, randomly sampled from the same dataset as in Sec. IV-A. Each image displayed a random target marked with a circle of 18 pixels radius; the participant’s task was to tell the system how to accurately place a point, marked with another circle of 5 pixels radius, on the designated target using a maximum of 5 prompt steps. Each participant received minimal instructions and, notably, was not informed about the existence of the correction module.
In the second phase, we provided the same participants with an explanation of the system’s functionalities, including two examples of prediction sentences and two correction sentences. With this new knowledge, they instructed the system to identify $10$ targets on $10$ different images.
We measured the distance between the predicted point and the target across prompts, as the distance between the centers of the two circles. The results (Fig. 8) demonstrate a significant improvement in task performance in both phases: all the users quickly learned how to use the system, and were able to bring the point close to the target with few corrections. Their performance was, as expected, better after the prompt expert suggestions (the median distance at the 5th prompt is 21 pixels, which is comparable with the target circle radius).
The quick and accurate target reaching demonstrates the effectiveness of the prediction-correction mechanism in Words2Contact for precise contact identification, essential for real robot applications. Participants also reported high engagement and satisfaction with the system.
@@ V Conclusions & Future Work
Words2Contact is especially useful in scenarios where users cannot directly control the robot, such as in human-robot collaboration settings and teleoperation. For instance, in a collaborative task, the robot could autonomously choose and adjust its hand placement on a shared work surface while the human partner focuses on a different aspect of the task. While the system adapts well in these contexts, the confirmation of the final contact point remains an open question. In this work, we mainly focused on language-based teleoperation, where the operator is remote from the scene and can confirm the location of the final contact point through visual feedback; however an equivalent confirmation system will also be necessary in human-robot collaboration. In our pilot study, users quickly adapted to the system, and real-world experiments demonstrated that Words2Contact delivers satisfactory contact placements, even in challenging environments, through its iterative correction mechanism. Based on these findings, we expect similar performance in human-robot collaboration, where a reliable confirmation system could ensure precise contact point adjustments during shared tasks.
For future work, we aim to reduce user reliance on visual feedback by enabling online corrections and scene-grounded trajectory generation for the end-effectors [shi2024yell, cui2023onlinelanguagecorrections, kwon2024language]. Additionally, improving the system’s performance will involve ensuring motion feasibility through real-time evaluation of inverse kinematics and surface safety using VLMs. For example, predicted contact points will be constrained to areas that are both a) reachable and b) safe, such as avoiding fragile surfaces like glass.
Finally, Words2Contact opens up opportunities for shared or full autonomy by capturing human-robot interaction data during teleoperation to build a robust dataset. This data could be used to fine-tune LLMs or train imitation learning approaches for shared autonomy [rouxel2024flowmatchingimitationlearning], allowing the system to autonomously manage routine tasks while requiring human intervention only for more complex decisions.